OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}(\frac{ 1^x + 2^x + 3 ^x ........n^x }{ n })^{\frac{ 1 }{ x }}\]

OpenStudy (anonymous):

Power is For Both nume and Deno

hartnn (hartnn):

can be solved by taking log on both sides...

hartnn (hartnn):

and also without LH, do you know what lim x->0 [a^x-1 /x] =... ?

hartnn (hartnn):

*[a^x-1]/x

OpenStudy (anonymous):

ln a/b

hartnn (hartnn):

\[\lim_{x \rightarrow 0}(\frac{ 1^x + 2^x + 3 ^x ........n^x }{ n })^{\frac{ 1 }{ x }}\\=\lim_{x \rightarrow 0}(\frac{ 1^x + 2^x + 3 ^x ........n^x +n-n}{ n })^{\frac{ 1 }{ x }} \\ \lim_{x \rightarrow 0}(\frac{ 1^x + 2^x + 3 ^x ........n^x+n }{ n }-1)^{\frac{ 1 }{ x }} \\\lim_{x \rightarrow 0}(\frac{ 1^x-1 + 2^x -1+ 3 ^x-1 ........n^x-1 }{ n }-1)^{\frac{ 1 }{ x }}\] got the hint ?

hartnn (hartnn):

and you need to take log on both sides first....to get 'x' in the denominator.

hartnn (hartnn):

am i on right track? @experimentX

OpenStudy (experimentx):

seems so ... ;let me do on copy!!

OpenStudy (experimentx):

oh .. that last one is +1

hartnn (hartnn):

yep, typo...

hartnn (hartnn):

\[\lim_{x \rightarrow 0}(\frac{ 1^x + 2^x + 3 ^x ........n^x }{ n })^{\frac{ 1 }{ x }}\\=\lim_{x \rightarrow 0}(\frac{ 1^x + 2^x + 3 ^x ........n^x +n-n}{ n })^{\frac{ 1 }{ x }} \\ \lim_{x \rightarrow 0}(\frac{ 1^x + 2^x + 3 ^x ........n^x-n }{ n }+1)^{\frac{ 1 }{ x }} \\\lim_{x \rightarrow 0}(\frac{ 1^x-1 + 2^x -1+ 3 ^x-1 ........n^x-1 }{ n }+1)^{\frac{ 1 }{ x }}\]

OpenStudy (experimentx):

although ... L'hopital seems a bit faster!!

hartnn (hartnn):

LH after taking log, right ?

OpenStudy (experimentx):

yeah!! your method works just fine as well!!

hartnn (hartnn):

hmm...much faster! thanks~

OpenStudy (experimentx):

oh!! i had been using unnecessary approximations. your method is faster!!

OpenStudy (experimentx):

should be n! e^(1/n)

OpenStudy (experimentx):

woops!! sorry (n!)^(1/n)

hartnn (hartnn):

using the standard formula for lim x->0 (1+x)^(1/x) and adjusting the exponents .

OpenStudy (experimentx):

the other would be use the usual unusual trick!! x = e^ln(x)

hartnn (hartnn):

haven't practised that much...would like to see steps, if you don't mind...

hartnn (hartnn):

why do i feel i have asked this limit Q..... O.o, @Yahoo! where are you ?

OpenStudy (anonymous):

I am Here..:)

OpenStudy (anonymous):

i also..Need to know the steps Plzz

OpenStudy (experimentx):